# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v1

@@ Abstract
Growing awareness of wellness has prompted people to consider whether their dietary patterns align with their health and fitness goals. In response, researchers have introduced various wearable dietary monitoring systems and dietary assessment approaches. However, these solutions are either limited to identifying foods with simple ingredients or insufficient in providing analysis of individual dietary behaviors with domain-specific knowledge. In this paper, we present DietGlance, a system that automatically monitors dietary in daily routines and delivers personalized analysis from knowledge sources. DietGlance first detects ingestive episodes from multimodal inputs using eyeglasses, capturing privacy-preserving meal images of various dishes being consumed. Based on the inferred food items and consumed quantities from these images, DietGlance further provides nutritional analysis and personalized dietary suggestions, empowered by the retrieval augmentation generation module on a reliable nutrition library. A short-term user study (N=33) and a four-week longitudinal study (N=16) demonstrate the usability and effectiveness of DietGlance.
@@ 1 Introduction
Dietary behaviors and human health are inextricably linked [72]. For instance, excessive consumption of processed foods high in sugars, unhealthy fats, and sodium increases the risk of chronic diseases [50]. Even small dietary improvements can significantly benefit health [28]. Thus, many individuals have sought dietary monitoring and analysis methods to refine their dietary choices. By offering insights into nutritional intake, these practices enable individuals to make informed decisions on diet planning, correct nutritional imbalances, and achieve health goals such as weight management or disease prevention [2]. The dominant dietary monitoring and analysis method nowadays is food journaling (on paper, websites, or mobile apps [69]), such as MyFitnessPal [53] and Boohee [10]. These tools allow users to log their food consumption, track nutrients, reflect on their dietary habits, and seek dietary recommendations from nutrition experts. However, these methods face several limitations.
@@ 2.1 Ingestive Episode Detection
Ingestive episode detection is essential for automatic food journaling [80]. Various wearable sensors have been explored for automated tracking. Wrist-mounted devices can track hand movements during eating [19, 33, 76], but these can be distracted by other gestures and require use on the dominant hand. More reliable approaches involve head- or neck-mounted devices like necklace [15, 86], headbands [6], caps [88], and ear-mounted sensors [5, 87, 7], though they are often uncomfortable and socially intrusive. In contrast, eyeglasses offer a more acceptable alternative, with their proximity to the mouth allowing for better detection of eating episodes. Previous work using eyeglasses includes sensors such as electromyography [85], piezoelectric [22], load cells [16], and acoustic sensors [46] for detection of ingestive behavior. Rahman et al. [60] utilized the Inertial Measurement Unit (IMU) in Google Glass for detection. IMU sensors combine accelerometers, gyroscopes, and occasionally magnetometers to measure the specific force, angular rate, and sometimes magnetic field of objects, enabling the tracking of motion and orientation in three-dimensional space [4]. Systems like FitByte [4] integrated IMU and proximity sensors, while MyDJ [69] utilized piezoelectric and accelerometer sensors.
@@ 2.3 Personalized Dietary Suggestions
Recent studies have explored machine learning models for personalized health and disease management suggestions. For example, Mitchell et al. [25] developed GlucoGoalie, which combines machine learning and a rule-based expert system to generate dietary goals for individuals with type 2 diabetes. To enhance flexibility and generalization, researchers have turned to LLMs for dietary guidance, leveraging their extensive knowledge and reasoning capabilities. Chatelan et al. [13] investigated the ability of ChatGPT to provide nutritional and dietary guidance for patients with type 2 diabetes and hemodialysis, and their findings show that the generated diet recipes are in accordance with the Diabetes Plate Method. Further, Yang et al. [83] introduced ChatDiet, an LLM-powered framework for personalized nutrition-oriented food recommender chatbots, and evaluated it on a case study including one-person information, where the generated food recommendation dialogue demonstrates the explainability and personalization. However, researchers noted that LLMs are easily struggling with hallucinations. One of study in [73] conducted an analysis involving registered dietitians to assess the capabilities of LLMs in providing nutritionally accurate and personalized information, showing the outputs of some instances do not align with standards upheld by registered dietitians and contain falsehoods, which can be misleading. Similarly, in [27], researchers found that while the nutritional advice proposed by ChatGPT is generally accurate, it has the potential to produce unsafe diets containing allergens.
@@ 4 System Design and Implementation
Based on the design requirements, we developed DietGlance, as shown in Fig. 2. First, we propose a multimodal sensing approach using camera, IMU, and audio sensors on eyeglasses to detect ingestion behaviors (Sec. 4.1) and capture meal images (Sec. 4.2) (R1, R4). The processed images and user profiles were analyzed using GPT-4V to identify the type and amount of food consumed (Sec. 4.3, R1). Unlike previous studies relying solely on LLM or requiring expert input, we further grounded the LLM with a reliable nutrition library through the RAG module to obtain nutritional analysis and personalized dietary suggestions (Sec. 4.3, R2, R3). Finally, we developed a novel user interface, distinct from previous studies, that allows users to interactively review and correct data, add details, explore results, and obtain more information (Sec. 4.4, R1-R4). All prompts are available in Appendix A.
@@ 4.1.2 Ingestive Episode Detection
After feature extraction, we employed a multimodal learning model (Fig. 4), building upon the framework introduced in [51], to classify time windows as ingestive or non-ingestive activities (R1). Specifically, instead of distinguishing between eating and drinking, we categorized both under the broader term “ingestive behaviors.” This decision accounts for practical challenges, such as the overlap in actions like consuming noodle soup (involving both eating and drinking) or swallowing soft foods without chewing. This simplification aligns with the study’s goals and improves usability in real-world applications.
@@ 4.2.2 Food, Drink, Tableware Segmentation
We further segmented the food, drinks, and tableware within the images. To exclude the food from nearby diners and include the user’s meal more completely in the egocentric images, we set the bottom one-third region as the primary area of interest (more illustration is available in the supplementary material 1.1). For segmentation, we employed Grounded-Segment-Anything (Grounded-SAM) [63], which enables precise identification based on textual descriptions. This approach allows the segmentation of a wide range of foods and drinks using class prompts and can easily accommodate new food types by adding new prompts. As illustrated in Fig. 5, we first utilized the pre-trained GroundingDINO model to detect objects based on the class prompt, then applied the pre-trained SAM model to convert detections into masks. Additionally, the class prompt (detailed in Table A1) included not only food and drink categories but also tableware, as these are often the only remaining objects post-meal.
@@ 4.3.1 Diet Identification
LLMs, trained on diverse datasets, excel at identifying various cuisines and ingredients in images and estimating food quantities through visual cues in a zero-shot manner, offering detailed nutritional information. In this study, we select GPT-4V among all LLMs due to its superior performance in diet recognition and estimation in open-world settings [42] (prompt available in Table A2). As shown in Fig. 6 (pink part), the meal log images and user profile (including gender, age, height, and weight) were fed into the GPT-4V to obtain the types and amount of each food and drink consumed by the user in this meal session.
@@ 4.3.2 RAG Module Construction
To build the RAG module, we have constructed a comprehensive nutrition library consisting of documents from a diverse collection of reliable resources. The inclusion criteria included (1) the government-approved food nutrient database, (2) official reports and booklets on food consumption and dietary suggestions for the general public and specific groups, issued by the government or the World Health Organization (WHO), (3) webpages from official government websites, containing recipes, scientific articles about food and nutrition, and healthy diet suggestions, (4) articles published by highly trustworthy organizations (e.g., National Institutes of Health), (5) peer-reviewed research articles from prestigious journals with 28 different needs (e.g., weight management, muscle development, and blood glucose stabilization). Detailed information is available in the supplementary material 1.2.
We then built the RAG module using LangChain framework [34]. The documents in the nutrition library were first split into chunks with LangChain’s RecursiveCharacterTextSplitter, resulting in 16,438 pages, as shown in Fig. 6 (blue part). Then OpenAI’s embedding model text-embedding-ada-002 created embeddings stored with Facebook AI Similarity Search (Faiss) [20] vector database. Based on these embeddings, the most relevant chunks will be searched based on the query to provide context for answering the query, enhancing the relevance and accuracy of the responses by incorporating the contexts. By invoking the retrieval chain created by LangChain, it would not only return the answer based on the prompt and query but also the sources of the answer from the relevant context, making the sources of dietary suggestions more transparent.
@@ 4.3.3 Nutritional Analysis and Dietary Suggestion
With the RAG-grounded LLM, we first use a crafted prompt (available in Table A3) to derive nutritional analysis based on the description and amount of food and drinks, including the nutrient and corresponding values. The total nutrition consumed per meal was calculated from the estimated portions. The system then checked if each nutrient intake was too high, too low, or within a reasonable range, offering suggested reference values.
@@ 4.4 Interface Design
To support users exploring the results, seeking further details, such as explanations for specific nutrients, alternative food options, or personalized meal plans, and providing contextual information, we developed an interactive desktop and mobile interface (Fig. 7) that not only displays results and suggestions but also allows users to correct errors, add missing details, and access more information. The interface consists of four major components: (1) Meal Image Logging (Fig. 7-A), (2) Interactive Nutritional Analysis (Fig.7-B), (3) Dietary Suggestion (Fig. 7-C), and (4) Context-Aware AI Chatbot (Fig. 7-D). Chronologically displayed meal log thumbnails allowed users to review their dietary patterns and access privacy-protected meal images (Fig. 7-A, R1, R4), while clicking on a meal session provided detailed nutritional analysis and total nutrient consumption (Fig.7-B, R1, R2). The interactive analysis enabled users to edit meal summaries and dynamically update nutritional insights and suggestions. Personalized dietary suggestions (Fig.7-C, R3), limited to seven items for clarity, offered general guidance for a balanced diet and tailored suggestions based on user goals, with transparent source references to enhance trust. The AI chatbot (Fig.7-D, R3) provided context-aware guidance, leveraging user profiles, meal logs, and prior interactions to deliver accurate, personalized responses, further streamlining the experience with preloaded common questions. This combination of features supported reflective, informed, and personalized dietary management. More details are available in the supplementary material 1.3.
@@ 5.1.1 Participants and Data Collection
We recruited 33 participants (P1-P33), including 21 self-declared males (63.6%) and 12 self-declared females (36.4%), aged 23 to 73 ($M=30.73$, $SD=10.51$). The group consisted of twenty-one university students, nine senior researchers, one accountant, one retired person, and one engineer. Twenty-one (63.6%) participants self-reported exp. with dietary monitoring and analysis, while the remaining participants (36.4%) were inexp. More details are available in Table A7.
Participants used Aria Glasses before, during, and after their meal sessions, capturing the ground truth video recordings, audio, and IMU data. Due to the power limitation of Aria Glasses, we did not ask the participants to wear the Aria Glasses all day but for periods that covered the meal sessions, which also included various behaviors like walking, sitting, talking, using mobile phones, etc. Participants reviewed their recordings and labeled the start and end of each ingestive episode, which were verified by the authors. Participants earned an average wage of about \$8 per hour of recording and would get about \$6.42 after finishing the user study additionally.
@@ 5.1.3 Study Settings
After obtaining participants’ consent, we introduced the system and provided a brief tutorial on the user interface. Participants were then asked to explore their dietary monitoring results, review meal identifications for accuracy, examine dietary suggestions and nutritional analysis, and use the chatbot. We recorded their screens and tracked interactions using Google Analytics [35]. Participants then completed a survey, including the System Usability Scale (SUS) [36], Net Promoter Score (NPS) [68], User Experience Questionnaire (UEQ) [67], and questions corresponding to the four design requirements. Finally, semi-structured interviews were conducted for open-ended feedback on their experience. Survey scales were converted to a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree) to make it consistent. Detailed scale items and interview scripts are in Table A9. The lead author first performed a deductive thematic analysis [11] using predefined themes derived from the system’s design requirements for the interviews and textual data from participants, and iteratively discussed with other authors to validate the codes and address any disagreements.
@@ 5.1.4 User Study Results
Twelve participants (36.4%) strongly agreed and 18 (54.5%) somewhat agreed that their privacy was secure, while three (9.1%) were neutral or disagreed. Similarly, 14 (42.4%) strongly and 16 (48.5%) somewhat agreed that data usage was transparent. Most participants trusted the system’s privacy protection due to clear communication and technical safeguards like data anonymization and secure storage. Exp. participants particularly appreciated the transparency of data handling (P15, exp.). Both groups valued visual privacy measures like background blurring, which “made everything except the food unrecognizable” (P9, inexp.). However, some concerns remained about external influences: “While I wasn’t worried, my friend sitting next to me was” (P27, inexp.). Exp. participants showed greater confidence and understanding of privacy protocols, while inexp. participants required more reassurance. These findings highlight the importance of improving privacy communication to address privacy concerns and enhance trust.
@@
Our evaluation revealed broadly positive usability, satisfaction, and user experience across both exp. and inexp. participants. Exp. participants consistently rated the system higher in terms of ease of use, automation, and efficiency, while inexp. participants highlighted challenges in adapting to workflows and interacting with the system. Satisfaction scores averaged 8.3 (NPS), with high utility acknowledged by most participants. However, the physical discomfort from prolonged wearing of glasses (such as heat buildup around the glasses’ legs, sweat accumulation, and slippage due to the limited glasses sizes provided) and workflow complexity emerged as areas for improvement, particularly for less tech-savvy users. Details on these assessments, including SUS, NPS, and UEQ metrics, are provided in the supplementary material 2.
@@
Participants spent an average of 10.10 minutes per session (2.59 minutes per meal). Reviewing dietary suggestions averaged 2.82 minutes. Inspecting nutritional analysis averaged 5.05 minutes, indicating participants prioritized understanding their diet over simply viewing suggestions. Manual corrections were minimal, with 27 changes across 144 meal sessions, mostly involving beverages (e.g., from “Coke” to “Coke Zero”). Users also corrected misidentified items or cooking methods (e.g., “grilled pork” to “grilled duck” or “steamed” to “fried”). Despite features designed to reduce errors of including nearby meals, occasional deletions of extra items indicate room for improvement. Participants also engaged actively with the chatbot, logging 125 conversation pairs (average of 3.79 per participant), primarily for meal planning with constraints such as “in Chinese style.” Iterative interactions, like P6 refining meal plans to exclude breakfast or avoid certain foods (full chat history is available in the supplementary material 3), demonstrated the chatbot’s adaptability in addressing preferences and fostering personalized dietary management. Overall, the system’s core features were actively utilized, with strong engagement enhancing the personalized user experience and demonstrating the system’s value in dietary monitoring and analysis.

### v2

@@ Abstract
Growing awareness of wellness has prompted people to consider whether their dietary patterns align with their health and fitness goals. In response, researchers have introduced various wearable dietary monitoring systems and dietary assessment approaches. However, these solutions are either limited to identifying foods with simple ingredients or insufficient in providing analysis of individual dietary behaviors with domain-specific knowledge. In this paper, we present DietGlance, a system that automatically monitors dietary in daily routines and delivers personalized analysis from knowledge sources. DietGlance first detects ingestive episodes from multimodal inputs using eyeglasses, capturing privacy-preserving meal images of various dishes being consumed. Based on the inferred food items and consumed quantities from these images, DietGlance further provides nutritional analysis and personalized dietary suggestions, empowered by the retrieval augmentation generation module on a reliable nutrition library. A short-term user study (N=33) and a four-week longitudinal study (N=16) demonstrate the usability and effectiveness of DietGlance, offering insights and implications for future AI-assisted dietary monitoring and personalized healthcare intervention systems using eyewear.
@@ 1 Introduction
Dietary behaviors and human health are inextricably linked [73, 47]. For instance, excessive consumption of processed foods high in sugars, unhealthy fats, and sodium increases the risk of chronic diseases [51]. Even small dietary improvements can significantly benefit health [27]. Thus, many individuals have sought dietary monitoring and analysis methods to refine their dietary choices. By offering insights into nutritional intake, these practices enable individuals to make informed decisions on diet planning, correct nutritional imbalances, and achieve health goals such as weight management or disease prevention [2]. The dominant dietary monitoring and analysis method nowadays is food journaling (on paper, websites, or mobile apps [70]), such as MyFitnessPal [54] and Boohee [11]. These tools allow users to log their food consumption, track nutrients, reflect on their dietary habits, and seek dietary recommendations from nutrition experts. However, these methods face several limitations.
@@
  • Through real-world implementation and studies, we contribute practical insights into the design and implementation of AI-assisted dietary monitoring and analysis systems using eyewear. These findings offer implications for future AI-assisted dietary monitoring and personalized healthcare intervention systems.
@@ 2.1 Ingestive Episode Detection
Ingestive episode detection is essential for automatic food journaling [76]. Various wearable sensors have been explored for automated tracking. In [28], Guan et al. developed a multi-source domain adaptation method to capture chewing activity for food type recognition utilizing IMU from wearable. Wrist-mounted devices can track hand movements during eating [19, 33, 78], but these can be distracted by other gestures and require use on the dominant hand. More reliable approaches involve head- or neck-mounted devices like necklace [87], headbands [7], caps [89], and ear-mounted sensors [6, 88, 8], though they are often uncomfortable and socially intrusive. In contrast, eyeglasses offer a more acceptable alternative, with their proximity to the mouth allowing for better detection of eating episodes. Previous work using eyeglasses includes sensors such as electromyography [86], piezoelectric [21], load cells [16], and acoustic sensors [46] for detection of ingestive behavior. Rahman et al. [61] utilized the Inertial Measurement Unit (IMU) in Google Glass for detection. IMU sensors combine accelerometers, gyroscopes, and occasionally magnetometers to measure the specific force, angular rate, and sometimes magnetic field of objects, enabling the tracking of motion and orientation in three-dimensional space [5]. Systems like FitByte [5] integrated IMU and proximity sensors, while MyDJ [70] utilized piezoelectric and accelerometer sensors.
@@ 2.3 Personalized Dietary Suggestions
Recent studies have explored machine learning models for personalized health and disease management suggestions. For example, Mitchell et al. [24] developed GlucoGoalie, which combines machine learning and a rule-based expert system to generate dietary goals for individuals with type 2 diabetes. To enhance flexibility and generalization, researchers have turned to LLMs for dietary guidance, leveraging their extensive knowledge and reasoning capabilities. Chatelan et al. [14] investigated the ability of ChatGPT to provide nutritional and dietary guidance for patients with type 2 diabetes and hemodialysis, and their findings show that the generated diet recipes are in accordance with the Diabetes Plate Method. Ataguba et al. [4] explored using LLM for personalized recipe generation and weight-loss management. Further, Yang et al. [84] introduced ChatDiet, an LLM-powered framework for personalized nutrition-oriented food recommender chatbots, and evaluated it on a case study including one-person information, where the generated food recommendation dialogue demonstrates the explainability and personalization. However, researchers noted that LLMs are easily struggling with hallucinations. One of study in [74] conducted an analysis involving registered dietitians to assess the capabilities of LLMs in providing nutritionally accurate and personalized information, showing the outputs of some instances do not align with standards upheld by registered dietitians and contain falsehoods, which can be misleading. Similarly, in [26], researchers found that while the nutritional advice proposed by ChatGPT is generally accurate, it has the potential to produce unsafe diets containing allergens.
@@ 4 System Design and Implementation
Based on the design requirements, we developed DietGlance, as shown in Fig. 2. First, we propose a multimodal sensing approach using camera, IMU, and audio sensors on eyeglasses to detect ingestion behaviors (Sec. 4.1) and capture meal images (Sec. 4.2) (R1, R4). The processed images and user profiles were analyzed using GPT-4V to identify the type and amount of food consumed (Sec. 4.3, R1). Unlike previous studies relying solely on LLM or requiring expert input, we further grounded the LLM with a reliable nutrition library through the RAG module to obtain nutritional analysis and personalized dietary suggestions (Sec. 4.3, R2, R3). Finally, we developed a novel user interface, distinct from previous studies, that allows users to interactively review and correct data, add details, explore results, and obtain more information (Sec. 4.4, R1-R4). All prompts are available in Supplement 1.
@@ 4.1.2 Ingestive Episode Detection
After feature extraction, we employed a multimodal learning model (Fig. 4), building upon the framework introduced in [52] and further integrating statistical features, to classify time windows as ingestive or non-ingestive activities (R1). Specifically, instead of distinguishing between eating and drinking, we categorized both under the broader term “ingestive behaviors.” This decision accounts for practical challenges, such as the overlap in actions like consuming noodle soup (involving both eating and drinking) or swallowing soft foods without chewing. This simplification aligns with the study’s goals and improves usability in real-world applications.
@@ 4.2.2 Food, Drink, Tableware Segmentation
We further segmented the food, drinks, and tableware within the images. To exclude the food from nearby diners and include the user’s meal more completely in the egocentric images, we set the bottom one-third region as the primary area of interest (detailed in Supplement 2.1). For segmentation, we employed Grounded-Segment-Anything (Grounded-SAM) [64], which enables precise identification based on textual descriptions. This approach allows the segmentation of a wide range of foods and drinks using class prompts and can easily accommodate new food types by adding new prompts. As illustrated in Fig. 5, we first utilized the pre-trained GroundingDINO model to detect objects based on the class prompt, then applied the pre-trained SAM model to convert detections into masks. Additionally, the class prompt (detailed in Supplement 1.1) included not only food and drink categories but also tableware, as these are often the only remaining objects post-meal.
@@ 4.3.1 Diet Identification
LLMs, trained on diverse datasets, excel at identifying various cuisines and ingredients in images and estimating food quantities through visual cues in a zero-shot manner, offering detailed nutritional information. In this study, we select GPT-4V among all LLMs due to its superior performance in diet recognition and estimation in open-world settings [42] (prompt available in Supplement 1.2). As shown in Fig. 6 (pink part), the meal log images and user profile (including gender, age, height, and weight) were fed into the GPT-4V to obtain the types and amount of each food and drink consumed by the user in this meal session.
@@ 4.3.2 RAG Module Construction
To build the RAG module, we have constructed a comprehensive nutrition library consisting of documents from a diverse collection of reliable resources. The inclusion criteria included (1) the government-approved food nutrient database, (2) official reports and booklets on food consumption and dietary suggestions for the general public and specific groups, issued by the government or the World Health Organization (WHO), (3) webpages from official government websites, containing recipes, scientific articles about food and nutrition, and healthy diet suggestions, (4) articles published by highly trustworthy organizations (e.g., National Institutes of Health), (5) peer-reviewed research articles from prestigious journals with 28 different needs (e.g., weight management, muscle development, and blood glucose stabilization). Detailed information is available in Supplement 2.2.
We then built the RAG module using LangChain framework [34]. The documents in the nutrition library were first split into chunks with LangChain’s RecursiveCharacterTextSplitter, resulting in 16,438 pages, as shown in Fig. 6 (blue part). Then OpenAI’s embedding model text-embedding-ada-002 created embeddings stored with Facebook AI Similarity Search (Faiss) vector database. Based on these embeddings, the most relevant chunks will be searched based on the query to provide context for answering the query, enhancing the relevance and accuracy of the responses by incorporating the contexts. By invoking the retrieval chain created by LangChain, it would not only return the answer based on the prompt and query but also the sources of the answer from the relevant context, making the sources of dietary suggestions more transparent.
@@ 4.3.3 Nutritional Analysis and Dietary Suggestion
With the RAG-grounded LLM, we first use a crafted prompt (available in Supplement 1.3) to derive nutritional analysis based on the description and amount of food and drinks, including the nutrient and corresponding values. The total nutrition consumed per meal was calculated from the estimated portions. The system then checked if each nutrient intake was too high, too low, or within a reasonable range, offering suggested reference values.
@@ 4.4 Interface Design
To support users exploring the results, seeking further details, such as explanations for specific nutrients, alternative food options, or personalized meal plans, and providing contextual information, we developed an interactive desktop and mobile interface (Fig. 7) that not only displays results and suggestions but also allows users to correct errors, add missing details, and access more information. The interface consists of four major components: (1) Meal Image Logging (Fig. 7-A), (2) Interactive Nutritional Analysis (Fig.7-B), (3) Dietary Suggestion (Fig. 7-C), and (4) Context-Aware AI Chatbot (Fig. 7-D). Chronologically displayed meal log thumbnails allowed users to review their dietary patterns and access privacy-protected meal images (Fig. 7-A, R1, R4), while clicking on a meal session provided detailed nutritional analysis and total nutrient consumption (Fig.7-B, R1, R2). The interactive analysis enabled users to edit meal summaries and dynamically update nutritional insights and suggestions. Personalized dietary suggestions (Fig.7-C, R3), limited to seven items for clarity, offered general guidance for a balanced diet and tailored suggestions based on user goals, with transparent source references to enhance trust. The AI chatbot (Fig.7-D, R3) provided context-aware guidance, leveraging user profiles, meal logs, and prior interactions to deliver accurate, personalized responses, further streamlining the experience with preloaded common questions. This combination of features supported reflective, informed, and personalized dietary management. More details are available in the Supplement 2.3.
@@ 5.1.1 Participants and Data Collection
We recruited 33 participants (P1-P33), including 21 self-declared males (63.6%) and 12 self-declared females (36.4%), aged 23 to 73 ($M=30.73$, $SD=10.51$). The group consisted of twenty-one university students, nine senior researchers, one accountant, one retired person, and one engineer. Twenty-one (63.6%) participants self-reported exp. with dietary monitoring and analysis, while the remaining participants (36.4%) were inexp. More details are available in Supplement 3.
Participants used Aria Glasses before, during, and after their meal sessions, capturing the ground truth video recordings, audio, and IMU data. Due to the power limitation of Aria Glasses, we did not ask the participants to wear the Aria Glasses all day but for periods that covered the meal sessions, which also included various behaviors like walking, sitting, talking, using mobile phones, etc. Participants reviewed their recordings and labeled the start and end of each ingestive episode, verified by the authors. Participants earned an average wage of about \$8 per hour of recording and would get about \$6.42 after finishing the user study additionally.
@@ 5.1.3 Study Settings
After obtaining participants’ consent, we introduced the system and provided a brief tutorial on the user interface. Participants were then asked to explore their dietary monitoring results, review meal identifications for accuracy, examine dietary suggestions and nutritional analysis, and use the chatbot. We recorded their screens and tracked interactions using Google Analytics [35]. Participants then completed a survey, including the System Usability Scale (SUS) [36], Net Promoter Score (NPS) [69], User Experience Questionnaire (UEQ) [68], and questions corresponding to the four design requirements. Finally, semi-structured interviews were conducted for open-ended feedback on their experience. Survey scales were converted to a 5-point Likert scale (1 = strongly disagree, 5 = strongly agree) to make it consistent. Detailed scale items and interview scripts are in Supplement 4. The lead author first performed a deductive thematic analysis [12] using predefined themes derived from the system’s design requirements for the interviews and textual data from participants, and iteratively discussed with other authors to validate the codes and address any disagreements.
@@ 5.1.4 User Study Results
Twelve participants (36.4%) strongly agreed and 18 (54.5%) somewhat agreed that their privacy was secure, while three (9.1%) were neutral or disagreed. Similarly, 14 (42.4%) strongly and 16 (48.5%) somewhat agreed that data usage was transparent. Most participants trusted the system’s privacy protection due to clear communication and technical safeguards like data anonymization and secure storage. Exp. participants particularly appreciated the transparency of data handling (P15, exp.). Both groups valued visual privacy measures like background blurring, which “made everything except the food unrecognizable” (P9, inexp.). However, some concerns remained about external influences: “While I wasn’t worried, my friend sitting next to me was” (P27, inexp.). “Even though the speech was unrecognizable, I was still reserved when discussing sensitive topics” (P26, exp.). In general, exp. participants showed greater confidence and understanding of privacy protocols, while inexp. participants required more reassurance. These findings highlight the importance of improving privacy communication to address privacy concerns and enhance trust.
@@
Our evaluation revealed broadly positive usability, satisfaction, and user experience across both exp. and inexp. participants. Exp. participants consistently rated the system higher in terms of ease of use, automation, and efficiency, while inexp. participants highlighted challenges in adapting to workflows and interacting with the system. Satisfaction scores averaged 8.3 (NPS), with high utility acknowledged by most participants. However, the physical discomfort from prolonged wearing of glasses (such as heat buildup around the glasses’ legs, sweat accumulation, and slippage due to the limited glasses sizes provided) and workflow complexity emerged as areas for improvement, particularly for less tech-savvy users. Details on these assessments, including SUS, NPS, and UEQ metrics, are provided in Supplement 5.
@@
Participants spent an average of 10.10 minutes per session (2.59 minutes per meal). Reviewing dietary suggestions averaged 2.82 minutes. Inspecting nutritional analysis averaged 5.05 minutes, indicating participants prioritized understanding their diet over simply viewing suggestions. Manual corrections were minimal, with 27 changes across 144 meal sessions, mostly involving beverages (e.g., from “Coke” to “Coke Zero”). Users also corrected misidentified items or cooking methods (e.g., “grilled pork” to “grilled duck” or “steamed” to “fried”). Despite features designed to reduce errors of including nearby meals, occasional deletions of extra items indicate room for improvement. Participants also engaged actively with the chatbot, logging 125 conversation pairs (average of 3.79 per participant), primarily for meal planning with constraints such as “in Chinese style.” Iterative interactions, like P6 refining meal plans to exclude breakfast or avoid certain foods (full chat history is available in Supplement 6), demonstrated the chatbot’s adaptability in addressing preferences and fostering personalized dietary management. Overall, the system’s core features were actively utilized, with strong engagement enhancing the personalized user experience and demonstrating the system’s value in dietary monitoring and analysis.